Order: Heparin Sulfate loading dose (bolus) 10,000 units IVP. Following the bolus, administer Heparin Sulfate continuous infusion 15 units/kg/hr.  

Available: Heparin 5,000 u/mL for the loading dose; Heparin 20,000 u/500 mL for the continuous infusion.

Patient weighs 110 pounds.

When setting the IV pump for the continuous infusion, how many mL/hour will the nurse administer in order to accomplish this order? Round your answer to the nearest whole number.

Please show me the work for this problem
Answer #1

ANSWER.

Calculation of ml/hr set for continuous infusion.

      Formula

            ml/hr = Volume (ml) /Time (hr).

    Amount (ml) required for continuous infusion.

          Formula = Desired dose /Dose in hand * Quantity.

                Here

                      Order = 15 units/kg/hr.

                  Weight of the patient = 110 pounds

               Convert pound to kg for calculation

                         1 kg = 2.205 pound.

              So

                  weight of the patient = 110 / 2.205

                                                   = 49.8

                                                   = 50 kg.

             Desired dose for patient with 50 kg (D) = 15 * 50

                                                                            = 750 units.

                 Dose in hand (H) = 20000 unit.

                                                    Quality (Q) = 500 ml.

          As per formula

              Amount (ml) required for continuous infusion = D/H *Q

                                                                    = 750 / 20000 * 500

                                                                    = 0.037 * 500

                                                                    = 18.5

                                                                    = 19 ml.

               ml/hr required for continuous infusion = Volume/Time

                                                                            = 19/1

                                                                            = 19 ml/hr.